Tunde will be 20yrs old in 3yrs, His sister will be 5yrs older in 2yrs. How old is Tunde’s sister now?
Found 2 solutions by rajagopalan, slegekp:
Answer by rajagopalan(174)

Tunde will be 20yrs old in 3yrs,
Tunde's present age=20-3=17 years
His sister will be 5yrs older in 2yrs.
So his sis will be 17+5 = 22 in 2 years
So present age of sis = 22-2=20.
How old is Tunde’s sister now?
Ans 20 years

Answer by slegekp(1)

20-3 = 17 (Tunde is 17yrs Old)
17+5 = 22 (Tunde's Present age and Tunde's sister age in 2yrs)
22-2 = 20 (Minus 2yrs older from 22 present)
Ans = 20